Circular No. 6869

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

S/1997 U 2 The recovery of the brighter of the new Uranian satellites is reported:

1998 UT α_{2000} δ_{2000} Observer
Mar. 27.79632 20^{h}57^{m}59.^{s}99 −17^{o}46′15′′.5 Smith
31.40380 20 58 34.18 −17 44 03.4 Hall

R. M. Smith, J. B. Jones and D. Windridge (Anglo-Australian Observa-

tory). 4-m reflector + CCD. Measurer B. Gladman. P. Hall and D. Graham (Cerro Tololo Interamerican Observatory). 4-m

reflector + CCD. Measurer J. J. Kavelaars.

These recovery positions are fully 30′′ from the prediction on IAUC 6834. Computations by G. V. Williams, K. Aksnes and the undersigned show that the reason for this is that there is another orbital solution from the 1984–1997 observations, much more strongly perturbed by the sun, separated from the sequence of solutions previously mentioned and corresponding to N ∼ 3.7. The 1998 recovery clearly shows that this high-e solution (see also IAUC 6780 and 6765) is the correct one.

Epoch = 1998 July 6.0 TT

T = 1997 June 16.0390 TT  ω = 18.^{o}0055
e = 0.509386  Ω = 255.8085 }2000.0
q = 0.040055 AU  i = 152.6686

a = 0.081643 AU n^{o} = 0.279383 P = 1289 days

1998 TT α_{2000} δ_{2000} Δ r_{U} ϵ Δα′′ Δδ′′
Apr. 7 20^{h}59.^{m}52 −17^{o}40.^{′}4 20.181 0.096 64.^{o}8 456+ 327+
17 21 00.72 −17 35.9 20.024 0.097 74.3 497+ 331+
27 21 01.61 −17 32.6 19.860 0.099 83.8 538+ 335+
May 7 21 02.18 −17 30.8 19.693 0.100 93.4 577+ 339+
17 21 02.41 −17 30.4 19.528 0.102 103.0 615+ 341+
27 21 02.31 −17 31.3 19.370 0.103 112.6 652+ 342+
June 6 21 01.89 −17 33.6 19.222 0.105 122.3 687+ 342+
16 21 01.17 −17 37.2 19.090 0.106 132.1 721+ 340+
26 21 00.18 −17 41.9 18.977 0.107 141.8 753+ 338+
July 6 20 58.96 −17 47.4 18.888 0.108 151.7 783+ 334+
16 20 57.56 −17 53.6 18.824 0.109 161.6 811+ 328+
26 20 56.04 −18 00.2 18.788 0.111 171.5 836+ 322+
Aug. 5 20 54.45 −18 07.0 18.781 0.112 178.4 860+ 314+
15 20 52.88 −18 13.6 18.804 0.113 168.6 881+ 306+
25 20 51.37 −18 19.7 18.855 0.114 158.6 900+ 297+
1998 April 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden